Optical Spectroscopy of the Niobium Methylidyne Free Radical, NbCH

\underline{David A. Gillett}, M. Barnes, G.F. Metha^{a}, K. Athanassenas and A.J. Merer, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B.C. V6T 1Z1, Canada.

Pulsed laser excitation spectra of niobium methylidyne (NbCH) have revealed 29 bands for NbCH and 17 for NbCD in the wavelength region 545–625 nm. NbCH was produced in a supersonic free jet expansion in helium by reacting methane with laser ablated niobium (355 nm Nd:YAG). The band structure can be organised into transitions involving excitation from the X̃^{3}Δ_{1} ground state to numerous vibronic levels of two close lying excited electronic states with Ω=2 and Ω=1, whose origins are near 16050 cm^{−1} and 16250 cm^{−1} respectively. Upper state assignments are complicated by a large vibronic (Renner-Teller) coupling between the two states and random perturbations due to ‘dark’ states. Wavelength resolved fluorescence spectra of NbCH(NbCD) have allowed us to determine two of the three X̃ state vibrational fundamentals:

ν_{2} = 603(466) cm^{−1}, ν_{3} = 905(858) cm^{−1}.

Work is currently underway to record the (0, 0) bands at cw laser resolution, which will allow accurate structural parameters to be determined and will resolve the niobium nuclear hyperfine structure. Modelling of the Renner-Teller interaction between the vibronic levels of the two close lying excited electronic states is also in progress.

aPresent address: School of Chemistry, University of Sydney, Building F11, Sydney, N.S.W. 2006, Australia